OpenStudy (anonymous):

Solve the following system of equations (using the Elimination method) and show all of your work. 2x + 5y = −3 −15x − 5y = −10

OpenStudy (texaschic101):

2x + 5y = - 3 -15x - 5y = - 10 -------------- - 13x = - 13 x = -13/-13 x = 1 now sub 1 into either of the original equations... 2x + 5y = - 3 2(1) + 5y = - 3 2 + 5y = - 3 5y = - 3 - 2 5y = - 5 y = - 1 check... -15x - 5y = - 10 -15(1) - 5(-1) = - 10 -15 + 5 = - 10 - 10 = - 10 (correct) ANSWER: x = 1 and y = - 1